OpenStudy (anonymous):

WILL GIVE METAL!!!! 18. A community group is planning the expansion of a square flower garden in a city park. If each side of the original garden is increased by 3 meters, the new total area of the garden will be 225 square meters. Find the length of each side of the original garden. 15 m 3 m 12 m √12 m

OpenStudy (anonymous):

it would be 12 because if you take 12 and add 3 (it says each side of the original garden increased by 3 meters) and then found the area of a square with side lengths of 15 (because 12 + 3 = 15), you would get 225.

OpenStudy (anonymous):

i just went through the answers and saw which ones worked and which ones didn't according to what the problem said

OpenStudy (anonymous):

can you possible help me with 3 more? If you cant then that's fine

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

Find the solutions to the system. y=x^2-2x-2 y=4x+5 *(-1,1) (-7,-23) *(-1,1) (7,33) *-1,33) (7,1) *No solution

OpenStudy (anonymous):

Solve the equation using the zero product property. (3x+1)(x-5)=0 *1/3 and 5 *-1/3 and -5 *-1/3 and 5 *1/3 and -5

OpenStudy (anonymous):

okay before i help, one quick question, do you factor or use the quadratic formula?

OpenStudy (anonymous):

Both but mostly factor

OpenStudy (anonymous):

what is the value of c so that y=x^2 +15x+c is a perfect square trinomial? *30 *15/2 *15/4 *225/4

OpenStudy (anonymous):

for the first one: okay so in order to do this you have to substitute. Since both of the equations are equal to "y", you can set them equal to each other. so, x^2-2x-2=4x+5 Then, what you want to do is put all the numbers onto one side and set the equation equal to zero so that you can factor.so what you wan tot do is subtract the 4x-5 from both sides, this leaves you with: x^2-6x-7=0 then, you factor and would get: (x-7) (x+1) So x= 7 or x= -1

OpenStudy (anonymous):

that's not an answer choice so would it be D:no solution or would it be C: (-1,33) and (7,1)

OpenStudy (anonymous):

Then, what you would have to do is plug the x's back into one of the original equations to get the solutions. so lets use y=4x+5 y= 4(7)+5 y= 28 +5 y=33 y+4(-1) + 5 y=1 so the answers are (7,33) and (-1,1)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

for the second one, I don't really remember the zero product property so i can't help you with that, sorry :( but for the last one, the answer is 30. not sure if you learned the formula for factoring a perfect square trinomial but its a^2 + 2ab + b^2 = (a+b)^2 so if you plug in 1,15, and 30 into that equation (a=1, b=15, and c=30) it would work

OpenStudy (anonymous):

wait. never mind i think i did that wrong

OpenStudy (anonymous):

ok thanks anyways. Possible help on one more its very similar to the first question asked.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Tasha is planning an expansion of a square flower garden in a city park. If each side of the original garden increased by 7m, the new total area of the garden will be 144m^2. Find the length of each side of the original garden. *19m *12m *5m *\[\sqrt{5}\]

OpenStudy (anonymous):

okay so same thing, i would plug in each answer to see if it works. i got 5m since if the original side is 5, and then you add 7 to it, you get twelve. Then, when you find the area of the shape, you would get 144